OpenStudy (anonymous):

How do I convert r^2 sin(2 theta) = 8 into Cartesian form?

zepdrix (zepdrix):

\[\Large\bf\sf r^2 \color{royalblue}{\sin(2 \theta)}\quad=\quad 8\]We want to mess with this blue part first. Do you remember your Sine Double Angle Identity? :)

zepdrix (zepdrix):

\(\Large\bf \color{#008353}{\text{Welcome to OpenStudy! :)}}\)

OpenStudy (anonymous):

Is that sin^2+cos^2=1? D:

zepdrix (zepdrix):

Nooo you silly billy. That's the Square Identity for Sine and Cosine. The Double Angle Formula for Sine is,\[\Large\bf\sf \sin(2\theta)=2\sin \theta \cos \theta\]Look familiar?

zepdrix (zepdrix):

\[\Large\bf\sf r^2 \cdot \color{royalblue}{2\sin \theta \cos \theta}\quad=\quad 8\]

OpenStudy (anonymous):

and where do I go from there? Do I divide 8 by 2sin(theta)cos(theta) and convert r^2 into x^2+y^2? Sorry so confused!!

OpenStudy (anonymous):

\[ r^22\sin\theta\cos\theta = 2(r\sin\theta)(r\cos\theta) \]